OpenStudy (anonymous):

please help. Before beginning a fundraiser, Lindsey estimated that the fundraiser would earn $500. She sold 114 raffle tickets at $5 each for the fundraiser, actually earning for the fundraiser. What was the percent error in Lindsey’s estimate? Round to the nearest tenth of a percent, if necessary. A. 12.3% B. 14% C. 70% D. 77.2% (medal)

OpenStudy (calculusxy):

Always use the formula that I gave you: \[\frac{ greater - lesser }{ actual } \times 100\]

OpenStudy (calculusxy):

The only "tricky" part here is that they just said "sold 114 raffle tickets at $5 each for the fundraiser, actually earning for the fundraiser." So that leaves you to find the actual price. This shouldn't be too hard because you just multiply: 114 times $5 and get your answer

OpenStudy (calculusxy):

@gknight

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

bro just do 114 times 5

OpenStudy (calculusxy):

1. 114 x 5 = ? 2. \[\frac{ $500 - ($14 \times 5) }{ ($14 \times 5) } \times 100\]

OpenStudy (calculusxy):

I have tried to make it as simple as I could.

OpenStudy (anonymous):

-614.2857 is this right @calculusxy

OpenStudy (anonymous):

i dont get it still

OpenStudy (anonymous):

Ok how many tickets did she sell times the price per ticket?

OpenStudy (anonymous):

57,000

OpenStudy (anonymous):

No, that's not it.

OpenStudy (anonymous):

570

OpenStudy (anonymous):

Ok that's it. Now what was the original prediction?

OpenStudy (anonymous):

$500

OpenStudy (anonymous):

Now the person before get how their equation works?

OpenStudy (anonymous):

what do you meen

OpenStudy (anonymous):

I'll show it. \[\frac{ 500 - (14 x 5) }{ (14 x 5) } x 100\] Make any sense? Remember PEMDAS?

OpenStudy (anonymous):

12.2807

OpenStudy (anonymous):

Hold on, I don't think thats right.

OpenStudy (anonymous):

Perenthesis first, its \[\frac{ 500 x 70 }{ 70 }\]

OpenStudy (anonymous):

500 times 70 is 35,000. What did you do next?

OpenStudy (anonymous):

fraction or divition

OpenStudy (anonymous):

I think it's divison. Correct? @calculusxy

OpenStudy (anonymous):

500

OpenStudy (anonymous):

Ok I re-did this. Do 112.3% 500 and it's about 561. So what you did was wrong. I'll try to explain this. You got the 114 being 570, the expected was 500. What is the difference?

OpenStudy (anonymous):

70

OpenStudy (anonymous):

Divide it by 5. What do you get?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

Ok, 500 is 100%, Add the 14% 114% What's 114% of 500?

OpenStudy (anonymous):

14%

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

No, 114%. Do 500 times 1.14 What is the result?

OpenStudy (anonymous):

570

OpenStudy (anonymous):

Ok then what is the answer to the question?

OpenStudy (anonymous):

C

OpenStudy (anonymous):

There you go. If you medal, medal the other guy.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Your welcome. Stay frosty,

OpenStudy (anonymous):

ok

OpenStudy (bella2112):

i understood nothing

OpenStudy (anonymous):

lol

OpenStudy (bella2112):

xD

OpenStudy (bella2112):

i kept reading it and i still didnt understand

OpenStudy (anonymous):

It's pretty easy to do, just gotta find the overestimate or underestimate difference. Multiple the 114, you get 570. Take the 570 and the 500. Find the difference. 70. The 70 divide by 5 is 14. As the 570 is above the original estimated it makes it a over percentage. Aka 114% On a calc it's 1.14.

OpenStudy (anonymous):

what do you mean by xD